OpenStudy (anonymous):

Solve for y: 1/y-5 =-2

OpenStudy (anonymous):

What do you think it is?

OpenStudy (trantom):

first add 5 1/y = 3 multiply both sides by y 1 = 3y Divide both sides by 3 1/3 = y Now check answer 1/(1/3)-5=-2 3-5 = -2 -2 = -2 The answer is correct; y=1/3